# MARS: A neurosymbolic approach for interpretable drug discovery

### v2

@@ 1 Introduction
Moreover, it is important to understand each drug’s mechanism-of-action (MoA), the molecular processes by which it achieves its medicinal effect. For instance, one drug, cortisone acetate, upregulates the activity of the glucocorticoid (GC) receptor protein, which, in turn, downregulates the cyclooxygenase (COX) protein. Since the COX protein is involved in creating inflammation, its inhibition reduces inflammation (Gonzalez-Cavazos et al., 2023). Thus, MoAs typically involve chains of physical, molecular interactions induced by a drug. Uncovering these interactions informs researchers as to how each drug works and de-risks potential side effects (Palve et al., 2021; Green et al., 2023).
@@
However, revealing MoAs alongside computational DD, a task we call MoA deconvolution, requires model interpretability: transparency into the processes or patterns which led to certain predictions (Molnar, 2022). Unfortunately, most state-of-the-art techniques on KGs rely on “black-box” models (Wu et al., 2020). Recently, neurosymbolic (NeSy) approaches, which combine logical rules with neural networks (DeLong et al., 2024) have been positioned as a promising avenue for MoA deconvolution. Specifically, NeSy methods tend to possess enhanced interpretability and better integrate structured domain knowledge, of which there is plenty in biomedicine (Acharya et al., 2023; Rigden & Fernández, 2023). However, since interpretability is broadly defined, there are no clear guidelines for assessing the plausibility of model interpretations, especially for this novel task (Molnar, 2022).
@@
Thus, we propose MoA deconvolution as a prediction task for evaluating interpretable methods on KGs. We generate a tailored KG for benchmarking this task, MoA-net. To perform MoA deconvolution, we introduce a NeSy DD approach called the MoA Retrieval System (MARS). To create MARS, we draw inspiration from previous NeSy methods (Liu et al., 2020; Drancé et al., 2021). However, unlike its predecessors, MARS leverages logical rules with learned rule weights. Following training, rule weights reflect relative usefulness to MARS’ reasoning processes.
@@
Alongside biomedical domain knowledge, MARS’ enhanced interpretability reveals a reasoning shortcut in which predictions are based on unintended semantics (Marconato et al., 2024b). Essentially, predictions upon MoA-net are driven by “degree-bias”, an artifact of node degree variance (Zietz et al., 2024), rather than the rules representing domain knowledge. Therefore, to address this second challenge, we consider the desiderata from Marconato et al. (2024a) for making NeSy systems shortcut-aware: (1) calibration, high accuracy on concepts unaffected by reasoning shortcuts, (2) performance, high accuracy despite reasoning shortcuts being present, and (3) cost effectiveness achieved through simple mitigation strategies. Using these desiderata as guidelines, we make MARS shortcut-aware for more insightful predictions involving DD and MoA deconvolution. Ultimately, our study underscores the importance of evaluating the capabilities of NeSy models within applied domains: by evaluating model interpretations against specific domain knowledge, we can more easily identify and mitigate shortcuts.
Contributions. In summary, our work:
  1. 1.
    devises a novel prediction task, MoA deconvolution, to assess the interpretability of NeSy and other KG-based approaches;
  2. 2.
    provides a novel biomedical KG, MoA-net, on which MoA deconvolution can be benchmarked;
  3. 3.
    presents the MoA retrieval system, MARS, a NeSy approach with enhanced interpretability through dynamic rule-weight updates;
  4. 4.
    through MARS’ enhanced interpretability, identifies a key reasoning shortcut affecting some NeSy models;
  5. 5.
    demonstrates ways to test and mitigate this reasoning shortcut, making MARS shortcut-aware and better aligned with domain knowledge.
@@ 2 Related Work and Background
Drug Discovery and MoA Deconvolution. Amongst previous open-source approaches for DD, many focus on drug repositioning (DR), as opposed to finding novel compounds (Rivas-Barragan et al., 2022; Urbina et al., 2021; Zheng et al., 2019). DR aims to discover new uses for drugs that have already been approved for distribution (Urbina et al., 2021). This prediction task allows researchers to leverage more available knowledge on approved drugs, including recorded clinical outcomes and side effects (Ratajczak et al., 2022; Liu et al., 2020). However, such relations are revealed during or after clinical trials, and are therefore rare or unknown for novel compounds. Furthermore, with such relations present, DR models oftentimes rely on associative patterns (e.g., drugs causing the same side effects also likely share indications) rather than mechanistic ones (e.g., involving physical, molecular interactions). Therefore, such models offer limited insight into drug MoAs.
NeSy AI for KGs. Several NeSy approaches for KGs involve logical rules in which the rule bodies reflect path-like patterns in the KG (e.g., $\texttt{treats}(A,D)\!\leftarrow\!\texttt{targets}(A,B)\land\texttt{regulates}(B,C)\land\texttt{attenuates}(C,D)$) (Liu et al., 2020; Dash & Goncalves, 2021; Sen et al., 2021). This is ideal for investigating MoAs, which typically occur as consecutive chains of molecular interactions (Crino, 2016). Some of these approaches utilize deep reinforcement learning (RL), in which a neural network contributes toward the optimization of a reward function (Acharya et al., 2023). In these specific cases, the reward function is partially determined by whether an agent’s trajectory through the KG corresponds to a rule body, thereby satisfying a rule. Therefore, the agent is guided by the path-like rules, and predictions are expected to align, to some extent, with such rules. In this study, however, we also identify a major risk: the agent will ignore rules if it can exploit other semantics for reward optimization. This results in reasoning shortcuts.
@@ 3 MARS: a NeSy approach for MoA deconvolution
Here, we build the MoA Retrieval System (MARS) [^1] https://github.com/laurendelong21/MARS to perform MoA deconvolution. MARS improves upon a method called Policy-guided walks with logical rules (PoLo) (Liu et al., 2020) by introducing dynamic, learned rule weights. This differs from previous approaches, where weights are static and pre-computed (e.g., mined or literature-derived) (Liu et al., 2020; Drancé et al., 2021). As discussed further, these learned weights make MARS shortcut-aware.
MARS takes two major inputs. The first is a KG. A KG uses nodes to represent entities and edges between them to represent relationships. A KG triple comprises two nodes connected by an edge of some specific type, or relation. For example, $\texttt{interacts}(Protein,Protein)$ states that two $Protein$ nodes are connected via the interacts relation. Node degree describes the number of edges connected to a node.
Specifically, the input KG must contain triples involving some relation of interest. To understand each MoA as the biological response to drug administration, we aim to investigate relations between drugs and biological processes (BPs), such as signal transduction or inflammation (Consortium, 2019) (i.e., $\texttt{induces}(Drug,BP)$). As per our knowledge, this is a novel application in the KG field. Further details on our KG are introduced within section 3.4. The second input includes metapaths of the KG with corresponding weights.
@@ 3.1 Metapath-based rules
Within this study, metapaths can be understood as a sequence of relations within the KG structure, making them inherently interpretable. In MARS, metapaths are used as the bodies of logical rules, in which the rule head, the left side of the implication arrow ($\Leftarrow$), constitutes a single, binary predicate representing the relation of interest between the first and last node types, e.g.,:
@@
For each metapath-based rule, $M_{i}$, in the set of feasible rules, =$\mathcal{M}=\{M_{1},M_{2},...,M_{m}\}$, we denote the rule weight by $w(M_{i})\in\mathbb{R}$, where $0\leq w(M_{i})\leq 1$. Such a weight indicates the relative usefulness of the metapath-based rule to the prediction task. In Section 3.3, we discuss how we initialize and compute these weights.
@@
Using a deep RL process, MARS trains an agent to take walks of length $L$ through the KG to connect pairs of nodes having the pre-defined relation of interest (e.g., induces); such edges are masked from the agent during training. Each walk generates a path, $P$, which can be understood as a series of $L$ transitions: $P:=(e_{c}\xrightarrow{r_{1}}e_{2}\xrightarrow{r_{2}}...\xrightarrow{r_{L}}e_{L+1})$. The agent may also remain at its current node. Ultimately, the goal of the agent is episodic: to find paths in which the starting node, $e_{c}$, and the terminal node, $e_{L+1}$, make up one of the true input pairs. By training the agent to do so, it can identify node pairs with some desired relationship, thus generalizing beyond the training set to predict novel pairs. Through a process akin to a Markov Decision Process (Bellman, 1957), the agent makes decisions about its next move based on information about its current position and the next possible actions. Additionally, however, the history of the agent’s previous actions are encoded with an LSTM (Hochreiter & Schmidhuber, 1997; Sherstinsky, 2020), whose parameters are trained to optimize the reward function, $R(S_{L+1})$ (Eq. 1), which is evaluated each time the agent completes $L$ transitions from some starting node, reaching a state $S_{L+1}$:
This reward function, adapted from Liu et al. (2020), quantifies how successful $P$ is according to two summands, where the hyperparameter $\lambda$ influences the balance between them. The first summand indicates whether the terminal node in the path, $e_{L+1}$, is one of the desired target (BP) nodes ($e_{d}$) that forms a true pair with the starting (drug) node, $e_{c}$. The second summand, contingent upon the first, indicates whether the corresponding metapath, $\tilde{P}$, matches any metapath-based rule, $M_{i}$. Therefore, the agent is not only encouraged to find connections between true pairs of nodes, but it is also guided toward paths which resemble known MoAs. Implementation details and hyperparameter selection are described in Appendices A.2 and A.4, respectively.
Of note, the second summand is proportional to some metapath-based rule weight, and MARS updates these weights during training. We accomplish this through a novel algorithm we call two-hop joint probability, or $P_{2H}$.
3.3 Weight updates via 2-hop joint probability
@@
After MARS is executed, its learned rule weights reflect each rule’s relative usefulness in the prediction task. Additionally, during training, weight updates drive the agent toward more informative paths and bypass the assumptions that pre-assigned rule weights are correct and informative. This eliminates the need for pre-computed or literature-derived rule weights; therefore, we initialize all weights uniformly as 0.5, a medium level of importance.
The naive way to implement weight updates ($\text{MARS}_{\text{naive}}$) is to increase weights according to the frequency at which each metapath-based rule is satisfied. However, this would not account for partial metapath matches. As an alternative, we propose updates via two-hop joint probability ($P_{2H}$). This metric approximates the usefulness of metapath-based rules based on the observed frequencies of the two-hop fragments that they comprise. Since the metapaths constituting rule bodies typically contain three or four consecutive relations, each two-hop fragment is extracted as in the following example:

### v3

@@ 1 Introduction
It is also important to understand each drug’s mechanism-of-action (MoA), the molecular processes by which it achieves its medicinal effect. For instance, as depicted in Fig. 1, MoAs typically involve chains or paths of physical, molecular interactions induced by a drug (Crino, 2016). Uncovering these interactions informs researchers as to how each drug works and de-risks potential side effects (Palve et al., 2021; Green et al., 2023). Revealing MoAs alongside computational DD, a task we call MoA deconvolution, requires model interpretability: transparency into the processes or patterns which led to certain predictions (Molnar, 2022). Unfortunately, most state-of-the-art techniques on KGs rely on “black-box” models (Wu et al., 2020). Recently, neurosymbolic (NeSy) approaches, which combine logical rules with neural networks (DeLong et al., 2024) have been positioned as a promising avenue for MoA deconvolution because they tend to possess enhanced interpretability.
Figure 1: MoA of cortisone acetate. Cortisone acetate upregulates the activity of the glucocorticoid (GC) receptor protein, which, in turn, downregulates the cyclooxygenase (COX) protein. Since COX is directly involved in creating inflammation, its inhibition reduces inflammation, thereby treating keratitis (Gonzalez-Cavazos et al., 2023). Data regarding protein interactions and biological processes (left) can be collected in a laboratory setting, whereas physiological effects like indications (right) are obtained during or after clinical trials. Created in https://BioRender.com.
However, interpretability is broadly defined (Molnar, 2022), which poses an additional challenge: there are no clear guidelines for assessing the plausibility of model interpretations, especially for this novel task. Although some previous studies present explainable or interpretable pipelines (Rivas-Barragan et al., 2022; Urbina et al., 2021), the corresponding explanations leverage associative patterns: two nodes with mutual connections are likely to share other connections (Paul et al., 2021). For example, such methods utilize associations regarding a drug’s pharmacological class (Ratajczak et al., 2022), side effects (Liu et al., 2021), or known indications (Fernández-Torras et al., 2022). Unfortunately, these associative patterns, which are discovered during or after clinical trials, are rare or absent for novel compounds. Furthermore, such patterns can not represent the MoA of a drug; instead, an MoA involves mechanistic patterns, such as the physical, molecular interactions shown in Fig. 1 (Gonzalez-Cavazos et al., 2023). Therefore, within this study, we focus upon model interpretability which provides mechanistic insight into drug MoAs.
@@
We propose MoA deconvolution, the prediction of mechanistic paths between drugs and their biological effects, as a prediction task for evaluating interpretable methods on KGs. To benchmark this task, we generate a tailored KG, called MoA-net, from real-world, experimental data. To perform MoA deconvolution, we introduce a NeSy DD approach called the MoA Retrieval System (MARS). To create MARS, we draw inspiration from previous NeSy methods (Liu et al., 2021; Drancé et al., 2021). However, unlike its predecessors, MARS achieves enhanced interpretability by learning weights associated with logical rules which resemble MoAs. Following training, rule weights reflect relative usefulness to MARS’ reasoning processes.
@@
Alongside biomedical domain knowledge, MARS’ enhanced interpretability reveals a reasoning shortcut in which predictions are based on unintended semantics (Marconato et al., 2024b). Essentially, predictions upon MoA-net are driven by “degree-bias”, an artifact of node degree variance (Zietz et al., 2024), rather than the rules representing domain knowledge. Therefore, to address this, we consider the desiderata from Marconato et al. (2024a) for making NeSy systems shortcut-aware: (1) calibration, high accuracy on concepts unaffected by reasoning shortcuts, (2) performance, high accuracy despite reasoning shortcuts being present, and (3) cost effectiveness achieved through simple mitigation strategies. Using these desiderata as guidelines, we make MARS shortcut-aware for more insightful predictions involving DD and MoA deconvolution. Ultimately, our study underscores the importance of evaluating the capabilities of NeSy models within applied domains: by evaluating model interpretations against specific domain knowledge, we can more easily identify and mitigate shortcuts.
@@ 2 Related Work and Background
NeSy AI for MoA deconvolution. Several NeSy approaches involve logical rules reflecting path-like patterns in biomedical KGs. For example, Sudhahar et al. (2024) investigates evidence chains, paths explaining associations between drugs and diseases. However, these explanations are derived separately from indication predictions, using an additional rule-mining model (Meilicke et al., 2019). Other approaches (Liu et al., 2021; Drancé et al., 2021) accomplish similar tasks through deep reinforcement learning (RL), in which a neural network contributes toward the optimization of a reward function (Acharya et al., 2023). In these specific cases, reasoning is guided by the path-like rules, and predictions are expected to align, to some extent, with such rules. In contrast to these previous studies, we focus upon paths representing MoAs, involving mechanistic, molecular relations. In this study, we also identify a major risk: the approach may neglect to utilize rules in favor of other semantics for reward optimization. This results in reasoning shortcuts.
@@ 3 MARS: a NeSy approach for MoA deconvolution
Here, we build the MoA Retrieval System (MARS) [^1] https://github.com/laurendelong21/MARS to perform MoA deconvolution. MARS improves upon a method called Policy-guided walks with logical rules (PoLo) (Liu et al., 2021) by introducing dynamic, learned rule weights. This differs from previous approaches, where weights are static and pre-computed (e.g., mined or literature-derived) (Liu et al., 2021; Drancé et al., 2021). As discussed further, these learned weights also make MARS shortcut-aware.
As shown in Fig. 2, MARS takes two major inputs. The first involves a KG. A KG uses nodes to represent entities and edges between them to represent relationships. A KG triple comprises two nodes connected by an edge of some specific type, or relation. Here, we represent triples as binary predicates: for example, the binary predicate $\texttt{interacts}(Protein,Protein)$ states that two $Protein$ nodes are connected via the interacts relation. Node degree describes the number of edges connected to a node.
Figure 2: Overview of the MoA retrieval system (MARS). Created in https://BioRender.com.
Specifically, the input KG must contain triples involving some relation of interest. As shown in Fig. 1, some information, such as indications, are discovered during or after clinical trials, so this information is typically unavailable for novel compounds. Therefore, to understand each MoA as the biological response to drug administration, we aim to investigate relations between drugs and biological processes (BPs), such as signal transduction or inflammation (Consortium, 2019) (i.e., $\texttt{induces}(Drug,BP)$). Specifically, we accomplish this through a link prediction task, in which we predict whether edges of type induces exist between $Drug$ and $BP$ nodes. Thereafter, new predictions regarding the induces relation serve as potential therapeutic outcomes for the chemical compound represented by the $Drug$ node. As per our knowledge, this is a novel application in the KG field. Further details on our KG are introduced within section 3.4. The second input, as depicted in Fig. 2, includes metapaths of the KG with corresponding weights.
@@ 3.1 Metapath-based rules
Within this study, metapaths can be understood as a sequence of triples within the KG structure, making them inherently interpretable. In MARS, metapaths are used as the bodies of logical rules, in which triples are connected by logical conjunctions ($\land$). Conjunctions indicate that, if all triples in the rule body are true, then the rule head is evaluated as true. The rule head, the left side of the implication arrow ($\Leftarrow$), is a single triple representing the relation of interest between the first and last node types of the metapath, e.g.,:
@@
If a rule head is evaluated as true, then the rule is satisfied. For each metapath-based rule, $M_{i}$, in a set of rules, =$\mathcal{M}=\{M_{1},M_{2},...,M_{m}\}$, we denote the rule weight by $w(M_{i})\in\mathbb{R}$, where $0\leq w(M_{i})\leq 1$. Such a weight indicates the relative usefulness of the metapath-based rule to the prediction task. In Section 3.3, we discuss how we initialize and compute these weights.
@@
[…]
Similarly to a Markov Decision Process (Bellman, 1957), the agent moves based on its current position and the next possible actions. Additionally, however, the history of the agent’s previous actions are encoded with an LSTM (Hochreiter & Schmidhuber, 1997; Sherstinsky, 2020), whose parameters are trained to optimize the reward function (Eq. 5 in Appendix A.4), which is evaluated each time the agent completes $L$ transitions from some starting node.
[…]
The second, supplementary reward, contingent upon the first, indicates whether the corresponding metapath, $\tilde{P}$, matches any metapath-based rule, $M_{i}$. The second reward is also proportional to the rule’s corresponding weight, and MARS updates these weights during training. We accomplish these updates through a novel algorithm we call two-hop joint probability, or $P_{2H}$. Therefore, the agent is not only encouraged to find connections between true pairs of nodes, but it is also guided toward paths which resemble known MoAs. Thus, MARS has two key interpretable features: (1) paths between nodes which serve as potential MoA predictions, and (2) learned rule weights which serve as a proxy for the importance of each rule.
3.3 MARS dynamically updates rule weights
@@
After MARS is executed, its learned rule weights reflect each rule’s relative usefulness in the prediction task. Additionally, during training, weight updates drive the agent toward more informative paths and bypass the assumptions that pre-assigned rule weights are correct. This eliminates the need for pre-computed or literature-derived rule weights; thus, we initialize all weights uniformly as 0.5, a medium level of importance. We test two approaches for updating rule weights.
Naive updates. The naive way to implement weight updates ($\text{MARS}_{\text{naive}}$) is to increase weights according to the frequency at which each metapath-based rule is satisfied. We record observed frequency, $O$, at which each metapath-based rule is satisfied. This is normalized by the batch-specific expected frequency, $E$, which assumes that every rule has a uniform probability across the total number of occurrences in that batch. If the agent finds zero occurrences, no weight updates are made. Ultimately, this produces a metric, $\mu$ (Eq. 1) in which $\mu>1$ indicates usefulness (the agent used that rule more than others), and $\mu<1$ indicates otherwise.
$\mu_{M_{i}}=O_{M_{i}}/E_{M_{i}}$ (1)
Notably, the value of $\mu$ is bound to avoid division by zero and extreme values. To adjust weight updates relative to batch size and rollouts, we define the minimum and maximum bounds on $\mu$ in Eqs. 2 and 3, where $\rho$ is the total number of metapath-based rules:
$\mu_{min}=\frac{\rho}{\text{batch size}\times\text{rollouts}}$ (2)
$\mu_{max}=\rho\times\text{batch size}\times\text{rollouts}$ (3)
Using Eq. 4, denoted by $\Phi$, $\mu$ (Eq. 1) is used to update the weight of a rule, $w(M_{i})$. Eq. 4 is regularized by the hyperparameter $\alpha\in\mathbb{R}$, where $0\leq\alpha\leq 1$ to control how subtle or drastic the weight update is, respectively. If $\alpha=0$, no weight updates are made. Therefore, $\alpha$ can be selected based on user needs or via hyperparameter optimization.
$\Phi(\mu,w(M_{i}))=w(M_{i})\times 2\alpha(\frac{\mu-1}{\mu+1})$ (4)
2-hop joint probability $P_{2H}$ updates. The second, more complex method to update weights is based on a term we coined, two-hop joint probability, $P_{2H}$. Pseudocode for $P_{2H}$ can be found in Algorithm 1 below. This metric approximates the usefulness of metapath-based rules based on full and partial matches. Since the metapaths constituting rule bodies contain several consecutive triples, each two-hop fragment is extracted as in the following example: